KhrisOnly123:

The table below shows the salvage value for raw steel. Tons of Steel Salvage Value 0 $0.00 2 $460.00 4 $920.00 6 $1,380.00 8 $1,840.00 10 $2,300.00 Based on the information in the table, what is the price per ton for raw steel? A. $230 B. $276 C. $184 D. $240

Krissy525:

Ok, so since the table goes by two tones, you would need to divide the cost of two tons by two. So, 460/2 is what?

velmalovesshaggy145:

its 210

KhrisOnly123:

210?

Krissy525:

No, it is not

velmalovesshaggy145:

i mean 230

velmalovesshaggy145:

missed the number

Krissy525:

Yes, but I was asking Khris, but Velma is right, so 230 is your answer

KhrisOnly123:

oh ok ty guys-

Krissy525:

No problem

velmalovesshaggy145:

@khrisonly123 wrote:
oh ok ty guys-
np

NoodlesAndRiceYT:

@velmalovesshaggy145, It is against the rules to give direct answers, please read the rules. https://questioncove.com/rules

1 attachment
velmalovesshaggy145:

oops my bad

cutebunny123:

You guys are right because if 2 tons cost 460 dollars, 460/2 = 230 the same is for all other alternatives 920/4 = 230 2300/10 = 230.

Krissy525:

That is true as well

velmalovesshaggy145:

Yes all three of us are correct